OpenStudy (anonymous):

help please integral of sqrt (x^2+6x) dx

OpenStudy (anonymous):

\[\int\limits \sqrt(x^2+6x) dx\]

OpenStudy (amistre64):

i dunno, there seems to be alot of variables to account for ....

OpenStudy (anonymous):

its really hard

myininaya (myininaya):

\[\int\limits_{}^{}\sqrt{x^2+6x+9-9} dx=\int\limits_{}^{}\sqrt{(x+3)^2-3^2} dx\] you will need a trig substitution

OpenStudy (anonymous):

(x^(2)+6x)*dx Multiply dx by each term inside the parentheses. dx^(3)+6dx^(2)

myininaya (myininaya):